# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v4

@@ 4.4 Nonce Distribution Security
Our threshold protocol produces nonces following an Irwin-Hall distribution rather than uniform. Using smooth Rényi divergence analysis [Raccoon2024], we prove this does not degrade security.
@@
Let $\Pi$ be our threshold ML-DSA with Irwin-Hall nonces. Then: $\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi^{*}}(\mathcal{A})}+q_{S}\cdot\epsilon$ where the per-coordinate smooth Rényi divergence satisfies $R_{2}^{\epsilon,\mathsf{coord}}\leq 1+n^{4}\beta^{2}/(4\gamma_{1}^{2})$ for $|S|=n$ signers. Since nonce coordinates are sampled independently, the full-vector divergence is $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ with $n\ell=1280$ for ML-DSA-65. The tail probability satisfies $\epsilon<10^{-30}$.
The square root arises from the smooth Rényi divergence probability transfer (Cauchy-Schwarz): for any event $E$, $\Pr_{P}[E]\leq\sqrt{R_{2}^{\epsilon}(P\|Q)\cdot\Pr_{Q}[E]}+\epsilon$. The $q_{S}\cdot\epsilon$ term accounts for the smoothing cost across $q_{S}$ signing queries. Since the nonce coordinates are sampled independently, $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the product of per-coordinate divergences. For ML-DSA-65 ($n\ell=1280$, $\mathsf{Adv}_{\Pi^{*}}\leq 2^{-192}$, $\epsilon<2^{-100}$):
$\mathsf{Adv}_{\Pi}\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot 2^{-192}}+q_{S}\cdot 2^{-100}$
$|S|$ $R_{2}^{\epsilon,\mathsf{coord}}$ $R_{2}^{\mathsf{vec}}$ $\sqrt{R_{2}^{\mathsf{vec}}}\cdot 2^{-96}$ Total bound
$q_{S}=2^{30}$ $q_{S}=2^{64}$
$\leq 17$ 1.003 $\approx 2^{5.4}$ $\approx 2^{-93.3}$ $\approx 2^{-70}$ $\approx 2^{-36}$
$\leq 25$ 1.014 $\approx 2^{25}$ $\approx 2^{-83.5}$ $\approx 2^{-70}$ $\approx 2^{-36}$
$\leq 33$ 1.041 $\approx 2^{75}$ $\approx 2^{-58.5}$ $\approx 2^{-58.5}$ $\approx 2^{-36}$
For $|S|\leq 17$, the $q_{S}\cdot\epsilon$ term dominates and the full-vector Rényi loss is only $\approx$3 bits. For $|S|\geq 25$, the $\sqrt{R_{2}^{\mathsf{vec}}}$ term becomes significant. This per-coordinate composition cost is inherent to the Rényi-based proof technique (shared with Raccoon [Raccoon2024]); tighter bounds via specialized Fiat-Shamir-with-Aborts reductions that exploit rejection sampling structure are an interesting open problem. Detailed calculations appear in Appendix D.
@@ 4.5 Malicious Security
Our protocol uses an optimistic approach: the normal signing path has no verification overhead. If failures exceed statistical expectation (e.g., >30>30 consecutive aborts, where honest probability is $(0.75)^{30}<10^{-4}$), a blame protocol identifies misbehaving parties (see Section 6.2).
@@ 4.6 Concrete Security Bounds
Share Privacy ($\mathbf{z}_{i}$) Statistical (min-entropy $\geq 24$K bits) N/A
@@
Table 3: Concrete security of masked threshold ML-DSA-65. Parameters match NIST Security Level 3. Underlying M-SIS hardness is 192 bits; the proven threshold bound depends on $|S|$ via the full-vector Rényi divergence $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$: for $|S|\leq 17$ and $q_{S}=2^{30}$, $\mathsf{Adv}<2^{-70}$ (see Theorem 4.11). Share privacy is statistical (no computational assumptions); see Theorem 4.4.
The threshold construction inherits the M-SIS/M-LWE hardness of single-signer ML-DSA. The proven forgery bound includes a $\sqrt{\cdot}$ loss from Rényi divergence and a $q_{S}\cdot\epsilon$ term from Irwin-Hall nonces (Theorem 4.11); tightening these bounds via specialized reductions is an open problem shared with Raccoon [Raccoon2024].
@@ 4.7.1 Profile P1 UC Security
The functionality $\mathcal{F}_{\mathsf{TEE}}$ operates as follows:
  • Round 2 Input: Receive masked commitments $\{(\mathbf{W}_{i},r_{i})\}_{i\in S}$ from parties (not broadcast)
@@
  • Round 3 Input: Receive response shares $\{(\mathbf{z}_{i},\mathbf{V}_{i})\}_{i\in S}$
  • Compute:
    1. 1.
      Aggregate $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}\bmod q$ and $c\mathbf{s}_{2}=\sum_{i}\mathbf{V}_{i}\bmod q$
    2. 2.
      Check z-bound: if $\|\mathbf{z}\|_{\infty}\geq\gamma_{1}-\beta$, output $\bot$
    3. 3.
      Check r0: if $\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},2\gamma_{2})\|_{\infty}\geq\gamma_{2}-\beta$, output $\bot$
    4. 4.
      Compute hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$
  • Final Output: Send $(\tilde{c},\mathbf{z},\mathbf{h})$ or $\bot$ to all parties; never reveal $c\mathbf{s}_{2}$ or individual $\mathbf{W}_{i}$
@@
Proof 4.14.
The proof is immediate by the ideal functionality abstraction. The simulator $\mathcal{S}$ operates as follows:
@@
    $\mathcal{S}$ simulates honest parties’ Round 1 messages (commitments $\mathsf{Com}_{h}$) and Round 3 messages (responses $\mathbf{z}_{h}$, masked shares $\mathbf{V}_{h}$). Responses have high min-entropy by Theorem 4.4 (statistical, no computational assumptions).
@@
  3. 3.
    $\mathcal{F}_{\mathsf{TEE}}$ computes the r0-check and hint internally, outputs only the final signature (or $\bot$).
Since $\mathcal{F}_{\mathsf{TEE}}$ never reveals $c\mathbf{s}_{2}$ or individual $\mathbf{W}_{i}$, the responses $\mathbf{z}_{h}$ have high min-entropy (Theorem 4.4, statistical), and the adversary’s view contains only $(\mathbf{w}_{1},\tilde{c},\sigma)$—all implicit in the signature—the real and ideal executions are indistinguishable. The distinguishing advantage is bounded by $\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$ (for the mask-based values only; the response privacy is statistical and contributes negligible statistical distance).
@@ 4.7.2 Profile P3+ Semi-Async Security Model
Definition 4.16 (Semi-Async Timing Model).
The semi-async model defines the following timing structure:
  • Pre-signing window $t_{0},t_{1}]$: Signers may precompute nonces $(y_{i},w_{i},\mathsf{Com}_{i})$ at any time
  • Challenge broadcast $t_{2}$: Combiner derives challenge $c$ from aggregated commitments
  • Response window $t_{2},t_{2}+\Delta]$: Signers respond within window $\Delta$ (e.g., 5 minutes)
  • 2PC execution $t_{3},t_{4}]$: Synchronous between CPs only
The adversary cannot control timing within honest parties’ windows but may observe message arrival times.
@@
Under the semi-async timing model, Protocol P3+ UC-realizes $\mathcal{F}_{\mathsf{ThreshSig}}$ against static adversaries corrupting up to T1T-1 signers and at most one CP, assuming:
@@
Security bound: $\epsilon\leq\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{OT}}+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
The proof appears in Appendix P. The key insight is that pre-signing security follows from commitment hiding (signers commit to nonces before seeing the challenge), and the response window does not introduce new attack vectors because:

### v5

@@ 4.4 Nonce Distribution Security
Our threshold protocol produces nonces following an Irwin-Hall distribution rather than uniform. Using smooth Rényi divergence analysis [Raccoon2024], we prove this introduces only a bounded security loss.
@@
Let $\Pi$ be our threshold ML-DSA with Irwin-Hall nonces. For any PPT adversary making $q_{S}$ signing queries:
$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\bigl(R_{2}^{\mathsf{vec}}\bigr)^{q_{S}/2}\cdot\sqrt{\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi^{*}}(\mathcal{A})}+q_{S}\cdot\epsilon$
where the per-coordinate smooth Rényi divergence satisfies $R_{2}^{\epsilon,\mathsf{coord}}\leq 1+|S|^{4}\beta^{2}/(4\gamma_{1}^{2})$ (see Appendix Q, Theorem Q.15; the formula uses the signing set size $|S|$, not the ring degree $n=256$). Since nonce coordinates are sampled independently, the full-vector divergence is $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ with $n\ell=1280$ for ML-DSA-65. The $(R_{2}^{\mathsf{vec}})^{q_{S}/2}$ factor accounts for multi-session composition (Remark Q.17). The tail probability satisfies $\epsilon<10^{-30}$.
Proof 4.18.
The smooth Rényi divergence probability transfer (Cauchy-Schwarz) gives: for any event $E$, $\Pr_{P}[E]\leq\sqrt{R_{2}^{\epsilon}(P\|Q)\cdot\Pr_{Q}[E]}+\epsilon$. For $q_{S}$ independent signing sessions, the joint Rényi divergence is $(R_{2}^{\mathsf{vec}})^{q_{S}}$ by the product rule (Definition Q.10), giving the $(R_{2}^{\mathsf{vec}})^{q_{S}/2}$ factor above. Since the nonce coordinates are sampled independently, $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$. For ML-DSA-65 ($n\ell=1280$, $\mathsf{Adv}_{\Pi^{*}}\leq 2^{-192}$, $\epsilon<2^{-100}$), the per-session cost is:
$q_{S}=1{:}\quad\mathsf{Adv}_{\Pi}\leq\bigl(R_{2}^{\mathsf{vec}}\bigr)^{1/2}\cdot 2^{-96}+2^{-100}$
$|S|$ $R_{2}^{\epsilon,\mathsf{coord}}$ $R_{2}^{\mathsf{vec}}$ Per-session bound ($q_{S}=1$): $\sqrt{R_{2}^{\mathsf{vec}}}\cdot 2^{-96}$
$\leq 17$ 1.003 $\leq 2^{5.5}$ ($\approx 2^{5.4}$) $\approx 2^{-93.3}$
$\leq 25$ 1.014 $\approx 2^{25}$ $\approx 2^{-83.5}$
$\leq 33$ 1.041 $\approx 2^{75}$ $\approx 2^{-58.5}$
Multi-session scaling. The conservative bound $(R_{2}^{\mathsf{vec}})^{q_{S}/2}$ grows rapidly: for $|S|=17$, it is vacuous for $q_{S}\gtrsim 36$ queries. The direct bound (Corollary Q.24) should always be used for concrete security estimates: the EUF-CMA loss is $0.013\cdot q_{S}$ bits for all $|S|\leq 33$ (Remark Q.17). Detailed conservative calculations appear in Appendix D.
@@ 4.5 Malicious Security
Our protocol uses an optimistic approach: the normal signing path has no verification overhead. If failures exceed statistical expectation (e.g., >33>33 consecutive aborts, where honest probability is $(0.75)^{33}<10^{-4}$), a blame protocol identifies misbehaving parties (see Section 6.1).
@@ 4.6 Concrete Security Bounds
Nonce Privacy ($\mathbf{y}_{i}$) Statistical ($\geq 20$K bits min-entropy) N/A
@@
Table 3: Concrete security of masked threshold ML-DSA-65. Parameters match NIST Security Level 3. Underlying M-SIS hardness is 192 bits; the proven threshold bound adds an Irwin-Hall nonce cost via $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$: per-session bound $\leq(R_{2}^{\mathsf{vec}})^{1/2}\cdot 2^{-96}\approx 2^{-93.3}$ for $|S|\leq 17$ (Theorem 4.17); over $q_{s}$ sessions the total loss is $0.013\cdot q_{s}$ bits (Corollary Q.24, Remark Q.17). The direct bound supports $q_{s}\leq 1{,}600$ with $\leq 10$-bit loss. Nonce privacy is statistical (no computational assumptions); key privacy given $\sigma$ follows computationally under M-LWE. See Theorem 4.7.
The threshold construction inherits the M-SIS/M-LWE hardness of single-signer ML-DSA. The proven forgery bound uses the direct shift-invariance analysis (Theorem Q.25, Corollary Q.24), giving <0.013<0.013 bits per signing session for $|S|\leq 33$ ($0.013\cdot q_{s}$ bits total; Corollary 4.3), resolving the scalability concern of prior work; developing an even tighter full-vector bound remains an interesting direction. The conservative Raccoon-style bound is in Remark 4.5.
@@ 4.7 Security Across Deployment Profiles
Definition 4.19 (Ideal Signing Functionality $\mathcal{F}_{\mathsf{ThreshSig}}$).
The functionality $\mathcal{F}_{\mathsf{ThreshSig}}(T,N,\mathsf{pp})$ operates as follows, where $\mathbf{z}_{i}\in R_{q}^{\ell}$ denotes each party’s response share and $\mathbf{V}_{i}\in R_{q}^{k}$ its r0-check share. On receiving $(\mathsf{SIGN},\mathsf{sid},\mu)$ from the combiner together with $T$ input shares $\{(\mathbf{z}_{i},\mathbf{V}_{i})\}_{i\in S}$ for $S\subseteq[N]$, $\mathcal{F}_{\mathsf{ThreshSig}}$:
  1. 1.
    Format check: If $|S|<T$ or any input share fails the format constraint ($\mathbf{z}_{i}\notin R_{q}^{\ell}$ or $\mathbf{V}_{i}\notin R_{q}^{k}$), output $\bot$ immediately to the combiner without querying $\mathcal{S}$.
  2. 2.
    Forwards $(\mathsf{SIGN},\mathsf{sid},\mu,S)$ to the simulator $\mathcal{S}$ and awaits a signature $\sigma$.
  3. 3.
    Outputs $\sigma$ to the combiner if $\mathsf{Verify}(\mathsf{pk},\mu,\sigma)=1$; otherwise outputs $\bot$.
Clarification on EUF-CMA vs. ideal functionality. The verification check in step 3 is a sanity check ensuring the simulator produces well-formed output; it does not implement EUF-CMA security (which is proven separately via game-based reduction). If an adversary constructs a valid forgery $(\mu^{\prime},\sigma^{\prime})$ for $\mu^{\prime}\neq\mu$, the ideal functionality accepts it in step 3, but this does not violate security—the adversary wins the EUF-CMA game, which occurs with negligible probability by Theorem 4.2.
The functionality reveals to $\mathcal{S}$: the message $\mu$, the signing-set $S$, and the outputs of corrupted parties ($\{(\mathbf{z}_{i},\mathbf{V}_{i})\}_{i\in C}$, where $C\subseteq S$ is the corrupted set). It does not reveal any honest party’s key share $(\mathbf{s}_{1,h},\mathbf{s}_{2,h})$ or nonce share $\mathbf{y}_{h}$ to $\mathcal{S}$.
@@ 4.7.1 Profile P1 UC Security
The functionality $\mathcal{F}_{\mathsf{TEE}}$ operates as follows.
@@
Round 3. Receive $\{(\mathbf{z}_{i},\mathbf{V}_{i})\}_{i\in S}$. Aggregate: $\mathbf{z}=\textstyle\sum_{i}\lambda_{i}\mathbf{z}_{i}\bmod q$, $c\mathbf{s}_{2}=\textstyle\sum_{i}\mathbf{V}_{i}\bmod q$. Abort with $\bot$ if $\|\mathbf{z}\|_{\infty}\geq\gamma_{1}-\beta$ or $\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},\,2\gamma_{2})\|_{\infty}\geq\gamma_{2}-\beta$. Otherwise compute $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\,\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},\,2\gamma_{2})$; output $(\tilde{c},\mathbf{z},\mathbf{h})$. Never reveal $c\mathbf{s}_{2}$ or individual $\mathbf{W}_{i}$.
@@
Proof 4.22.
We construct an explicit simulator $\mathcal{S}$ for each protocol round and prove indistinguishability.
Simulator construction.
@@
    Round 0 (DKG): $\mathcal{S}$ generates honest key shares $\mathbf{s}_{1,h},\mathbf{s}_{2,h}$ and nonce shares $\mathbf{y}_{h}$ inside the simulated $\mathcal{F}_{\mathsf{TEE}}$, following the real key DKG and nonce DKG protocols. Since DKG outputs to corrupted parties are determined by the $(T,N)$-Shamir sharing structure, $\mathcal{S}$ produces consistent evaluations $\{\mathbf{f}_{h}(j)\}_{j\in C}$.
@@
    Round 1 (Commitments): $\mathcal{S}$ samples $\tilde{\mathbf{y}}_{h},\tilde{\mathbf{w}}_{h},\tilde{r}_{h}$ from the simulated DKG, computes the committed value $\mathsf{Com}_{h}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$ uniformly, and programs the random oracle so that $H(\mathsf{``com^{\prime\prime}}\|\tilde{\mathbf{y}}_{h}\|\tilde{\mathbf{w}}_{h}\|\tilde{r}_{h})=\mathsf{Com}_{h}$ (using the same domain separator as the real commitment scheme $\mathsf{Com}_{i}=H(\mathsf{``com^{\prime\prime}}\|\mathbf{y}_{i}\|\mathbf{w}_{i}\|r_{i})$). By commitment hiding (preimage resistance of $H$), the uniform $\mathsf{Com}_{h}$ is indistinguishable from real commitments. If the blame protocol requests an opening of $\mathsf{Com}_{h}$, $\mathcal{S}$ provides $(\tilde{\mathbf{y}}_{h},\tilde{\mathbf{w}}_{h},\tilde{r}_{h})$, which verifies correctly by the programmed RO.
  3. 3.
@@
  4. 4.
    Round 3 (Responses): $\mathcal{S}$ computes $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ from simulated DKG values. This is identical to the real computation (SD =0=0). Masked shares $\mathbf{V}_{h}=\lambda_{h}c\mathbf{s}_{2,h}+\mathbf{m}_{h}^{(s2)}$ are computed analogously.
Indistinguishability.
The adversary’s view in the real and simulated executions is indistinguishable for the following reasons:
  • Response $\mathbf{z}_{h}$: Computed identically from simulated DKG values (SD =0=0). The simulator has access to the honest party’s key share $\mathbf{s}_{1,h}$ because it generated $\{\mathbf{s}_{1,j}\}$ during the simulated DKG (Round 0, setup step). The nonce share $\mathbf{y}_{h}$ has high conditional min-entropy by Theorem 4.7 (statistical); key privacy given $\sigma$ is conjectured to hold computationally under Module-LWE (open problem; see Remark 4.9).
  • TEE outputs: $\mathcal{F}_{\mathsf{TEE}}$ hides all internal state (individual $\mathbf{W}_{i}$, $c\mathbf{s}_{2}$); only $(\mathbf{w}_{1},\tilde{c},\sigma)$ exit the enclave, all of which are implicit in the signature.
  • Masks: Replacing honest-honest PRF terms with uniform introduces advantage $\leq\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
Abort handling.
The honest per-attempt success probability is $\approx 25\%$ (rejection sampling). After >33>33 consecutive failures, honest probability is $(0.75)^{33}<10^{-4}$, triggering the blame protocol (Section 6.1). The simulator handles aborts by forwarding $\bot$ from $\mathcal{F}_{\mathsf{TEE}}$.
Total distinguishing advantage: $\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$ (nonce share privacy is statistical and contributes zero to the PRF-based advantage).
@@ 4.7.2 Profile P3+ Semi-Async Security Model
Definition 4.24 (Semi-Async Timing Model (Formalized)).
The semi-async model extends the standard UC framework with bounded-time windows. Execution proceeds in optimistic synchronous phases with timeout-based fallback:
  1. 1.
    Pre-signing window $t_{0},t_{1}]$: Signers may precompute nonces $(\mathbf{y}_{i},\mathbf{w}_{i},\mathsf{Com}_{i})$ at any time before $t_{1}$. This is an asynchronous phase; early precomputation is allowed but not required.
  2. 2.
    Challenge broadcast $t_{2}$: Combiner derives challenge $c$ from aggregated commitments and broadcasts $(c,\mu)$ to all signers at time $t_{2}$. This is a synchronous event (all parties receive the broadcast at the same logical time in the UC model).
  3. 3.
    Response window $t_{2},t_{2}+\Delta]$: Honest signers should respond within window $\Delta$ (e.g., 5 minutes for human-in-the-loop scenarios). Adversary scheduling rules:
    • Honest party delivery: Messages from honest signers arriving within $t_{2},t_{2}+\Delta]$ are guaranteed to be delivered to the combiner.
    • Adversarial delay: The adversary may delay delivery of honest parties’ messages by up to $\Delta$, but cannot prevent delivery entirely (eventual delivery assumption).
    • Timeout: If fewer than $T$ responses arrive by $t_{2}+\Delta$, the combiner aborts the session and outputs $\bot$. This is not a security failure—it is an availability event (honest signers who responded late are still secure; no key material is leaked).
    • Message reordering: The adversary may reorder messages from different parties arriving within the window, but cannot inject or modify honest parties’ messages (authenticated channels assumption).
  4. 4.
    2PC execution $t_{3},t_{4}]$: Synchronous 2PC between the two CPs (standard UC model applies; this phase has no timing relaxation).
UC adversary capabilities. The adversary observes message arrival times (timestamps) within the response window but cannot control the content or the fact of delivery for honest parties. Replay attacks are prevented via unique session identifiers $\mathsf{sid}$ binding each response to $(c,\mu)$. The model is semi-asynchronous in the sense that signers do not need to coordinate their response times (unlike multi-round MPC), but the protocol remains UC-secure as long as $\geq T$ responses arrive within $\Delta$.
@@
Under the semi-async timing model, Protocol P3+ UC-realizes $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ in the $(\mathcal{F}_{\mathsf{OT}},\mathcal{F}_{\mathsf{GC}})$-hybrid model against static adversaries corrupting up to T1T-1 signers and at most one CP, assuming:
@@
Security bound: $\epsilon\leq\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{ext}}+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$, where $\epsilon_{\mathsf{ext}}\leq 2^{-B/2}$ is the LP07 cut-and-choose extraction error (for corrupted $\mathsf{CP}_{1}$). By UC composition (using the nonce DKG and signing sub-protocols as ideal functionalities), this implies that Protocol P3+ UC-realizes the full $\mathcal{F}_{\mathsf{ThreshSig}}$ under the same assumptions.
Proof 4.26 (Proof sketch (full proof in Appendix P)).
The simulator replaces OT and GC with ideal functionalities ($\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{GC}}$), then simulates pre-signing by sampling $\tilde{\mathbf{y}}_{i}$ and computing $\tilde{\mathbf{w}}_{i}=\mathbf{A}\tilde{\mathbf{y}}_{i}$ (Lemma P.8, SD =0=0), and simulates responses via $\mathbf{z}_{i}=\tilde{\mathbf{y}}_{i}+c\cdot\mathbf{s}_{1,i}$ (Lemma P.10, SD =0=0). The response window does not introduce new attack vectors because: